OpenStudy (anonymous):

A piece of circular bread has a diameter of 11 in. What is the approximate area of this piece of bread? Use 3.14 to approximate pi and round your answer to the nearest hundredth. _______in(2)

OpenStudy (anonymous):

@iambatman @Ilovecake @inowalst @TheSmartOne

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

plz help @TheSmartOne @inowalst @iambatman

OpenStudy (anonymous):

i think 47.5

OpenStudy (anonymous):

I heard someone needed help

OpenStudy (anonymous):

cumonn, do u mind? haha

OpenStudy (anonymous):

Lol Hold your horses... Dinner's coming

OpenStudy (anonymous):

Diameter is a whole so cut it in half that will be your radius

OpenStudy (anonymous):

u guys can just give me a straight answer at this point

OpenStudy (anonymous):

just plz hurry

OpenStudy (anonymous):

use the formula d=2*r 11=2*r 11/2=r r=5.5 or 6 A=3.14*r*r A=3.14*6*6 A=113.04

OpenStudy (anonymous):

113 in(2) is your answer

OpenStudy (anonymous):

94.99 in^2\[A=3.14*\left( 11/2 \right)^2\]

OpenStudy (anonymous):

^ is correct :/